# Larimar: Large Language Models with Episodic Memory Control

### v1

@@ Abstract
Efficient and accurate updating of knowledge stored in Large Language Models (LLMs) is one of the most pressing research challenges today. This paper presents Larimar - a novel, brain-inspired architecture for enhancing LLMs with a distributed episodic memory. Larimar’s memory allows for dynamic, one-shot updates of knowledge without the need for computationally expensive re-training or fine-tuning. Experimental results on multiple fact editing benchmarks demonstrate that Larimar attains accuracy comparable to most competitive baselines, even in the challenging sequential editing setup, but also excels in speed—yielding speed-ups of 4-10x depending on the base LLM —as well as flexibility due to the proposed architecture being simple, LLM-agnostic, and hence general. We further provide mechanisms for selective fact forgetting and input context length generalization with Larimar and show their effectiveness.
@@ 1 Introduction
Table 1: Comparison between different editing methods from the requirement and capability perspective. ROME and GRACE from existing editing methods are shown. The wall clock time for a single edit (averaged over 10 edits) on the CounterFact dataset for ROME [Meng et al., 2022a] and GRACE [Hartvigsen et al., 2022] was computed using the EasyEdit [Wang et al., 2023a] framework with a single A100 (80G) GPU.
@@ 2.2 Memory inference
Implementation:
We employed a BERT large encoder [Devlin et al., 2018] combined with either a GPT2-large [Radford et al., 2019] or a GPTJ-6B decoder and a memory matrix (512x768) for our training experiments, naming the resulting models Larimar-1.3B and Larimar-6B, respectively. Our training data comprised 7.6 million examples constructed by splitting WikiText  [Merity et al., 2016] texts to small chunks of 64 tokens. In testing, the Larimar-1.3B model achieved a perplexity of 14.6, while the Larimar-6B model reached 15.9 on 1,000 random WikiText samples, indicating that adding memory barely affects performance. We trained Larimar-6B models for 10 epochs using Adam optimizer, learning rate 5e-6 and batch size 32. For the Larimar-6B’s training, we used a setup with eight NVIDIA A100-80GB GPUs on a single node, utilizing bfloat16 precision and PyTorch Lightning with the DeepSpeed ZeRO Stage 2 for efficient distributed training.
@@ 3 Memory operations
1       // $\mathbf{Z}$ encoding of the episode to be written to memory (i.e. $\mathbf{Z}=e(\mathbf{X})$) Sample $\xi\sim\mathcal{N}(0,\sigma_{\xi}^{2}I)$   Let $\mathbf{Z}_{\xi}=\mathbf{Z}+\xi$   Compute addressing weight $\mathbf{W}_{0}=\mathbf{Z}_{\xi}\mathbf{M}_{0}^{\dagger}$     // $\mathbf{M}_{0}$ is a learned parameter representing prior memory Compute posterior memory $\mathbf{M}=\mathbf{W}_{0}^{\dagger}\mathbf{Z}_{\xi}$return $\mathbf{M}$
@@
4       // $\mathbf{M}$ posterior memory from previous write // $\mathbf{Z}$ encoding of the read input (ie. $\mathbf{Z}=e(\mathbf{X})$) Compute mean addressing weight $\overline{\mathbf{W}}=\mathbf{Z}\mathbf{M}^{\dagger}$   Sample $\mathbf{W}\sim\mathcal{N}(\overline{\mathbf{W}},\sigma_{\mathbf{W}}^{2}I)$   // $\sigma_{\mathbf{W}}$ is a learned parameter Compute output latent $\mathbf{Z}_{\text{read}}=\mathbf{WM}$return $\mathbf{Z}_{\text{read}}$
@@
7       // $\mathbf{M}$ is the posterior memory from a previous write Sample $\mathbf{W}\sim\mathcal{N}(0,I)$  Compute output latent $\mathbf{Z}=\mathbf{WM}$return $\mathbf{Z}$
@@ 5.1 Wall Clock time
Table 1 presents the wall clock time for each editing method across 10 edits, calculated within the EasyEdit framework [Yao et al., 2023] on a single A100 GPU. Results show that Larimar is 4-10x times faster compared to ROME [Meng et al., 2022a] and GRACE [Hartvigsen et al., 2022], two most competitive existing LLM editing baselines. Table 7 in Appendix further provides a edit-time comparison within other existing baselines, as shown in [Yao et al., 2023], establishing Larimar’s advantage on high-speed editing. Table 1 further lists Larimar’s abilities to handle edits in a training- or tracing- free manner, enabling high-speed editing, handling selective forgetting, and maintain ability to sequential editing setup.
Editor Edit Success Paraphrase Neighborhood S M S M S M GPT-2 XL 22.2 -4.8 24.7 -5.0 78.1 5.0 $*{}^{*}$ 100.0 98.8 87.9 46.6 40.4 -6.2 $*{}^{*}$ 99.1 91.5 48.7 28.9 70.3 3.5 KN 28.7 -3.4 28.0 -3.3 72.9 3.7 KE 84.3 33.9 75.4 14.6 30.9 -11.0 KE-CF 99.9 97.0 95.8 59.2 6.9 -63.2 MEND 99.1 70.9 65.4 12.2 37.9 -11.6 MEND-CF 100.0 99.2 97.0 65.6 5.5 -69.9 ROME 100.0 97.9 96.4 62.7 75.4 4.2 Larimar-1.3B 100.0 99.8 41.7 0.4 74.7 1.6 GPT-J 16.3 -7.2 18.6 -7.4 83.0 7.3 FT 100.0 99.9 96.6 71.0 10.3 -50.7 FT+L 99.6 95.0 47.9 30.4 78.6 6.8 MEND 97.4 71.5 53.6 11.0 53.9 -6.0 ROME 99.9 99.4 99.1 74.1 78.9 5.2 PROMPT 99.7 80.9 91.0 32.9 37.9 -2.8 IKE (w/ 32 demonstrations) 100.0 91.7 95.2 64.5 77.0 35.2 IKE (w/o paraphrases) 100.0 73.8 83.4 IKE (w/o neighbors) 100.0 99.8 11.5 Larimar-6B 99.6 96.0 76.5 22.4 80.2 3.9
Table 2: Single fact edit performance on CounterFact dataset comparing Larimar against baseline. Top two best systems are highlighted. Unlike other methods Larimar uses dynamic memory updates with memory-conditioned decoding, and does not require gradient update on edit samples, as opposed to methods that require training (FT, FT+L, MEND) or tracing plus decoder updating (ROME) on edit samples (ROME), or in-context demonstrations (IKE) of (paraphrased) edits and neighboring samples retrieved from a corpus.
@@ 5.2 Single Fact editing
We compare the performance of Larimar against a number of recently proposed knowledge editing approaches on the CounterFact dataset [Meng et al., 2022a] designed for testing language models handling of counterfactual edits. It includes 21,919 records to assess if the models can learn new facts rather than simply memorizing the target words. Following other works [Meng et al., 2022a, Zheng et al., 2023], we used the first 2000 samples of this dataset and report the average over single fact editing results for Larimar-1.3B and Larimar-6B in Table 2. The performance scores for baselines are from [Meng et al., 2022a, Zheng et al., 2023] (see Related Work and Appendix for details on baseline methods). As opposed to training the LLM on edits, or causally tracing the original fact within LLM and updating the relevant parameters to reflect edit, we leverage Larimar’s one-shot memory update for editing. Wherein, the memory posterior is updated as the edit(s) of interest is written, and then the updated memory is queried. The read-out from the memory then conditions the decoder to output the edit.
@@
As can be seen, when compared to existing editing baselines, Larimar achieves comparable performance in successfully editing new facts, and in the ability to handle neighborhood prompts (on par with ROME, when based on GPT-2XL, and better performance, when based on GPT-J), while there remains room to improve generalization. When compared to existing in-context editing approaches (PROMPT and IKE baselines) [Zheng et al., 2023], Larimar does not need multiple in-context demonstrations of the edit and its paraphrases, as well as of neighboring facts, to the decoder, which are retrieved from a corpus. However, as we show in the Appendix (Table 8), when Larimar has access to one additional paraphrase per fact, by writing it in the memory, the generalization performance increases from 76.5 to 82.8. Note that in this setup the average number of added paraphrase per fact is one and we queried the model with a paraphrased prompt unseen by the memory. Further, ablation experiments in Appendix shows that a scope detector, either trained on Larimar encodings or encodings from an external LLM, helps with better paraphrase generalization and neighborhood specificity. Throughout the paper, Larimar is configured with a scope detector, unless otherwise mentioned. For details, see Appendix.
We also evaluated Larimar on the ZsRE benchmark [Levy et al., 2017], a QA dataset for relation extraction through reading comprehension, with results displayed in Appendix in Table 12. Performance scores for GPT-2 XL based baselines are cited from [Meng et al., 2022a], whereas performance of ROME on GPT-J was independently estimated by us. Unlike the CounterFact evaluation, this assessment uses exact match counts for scoring $\mathbb{I}[o^{*}=\text{argmax}_{o}\mathbb{P}[o]]$. Compared to baselines, Larimar demonstrates effective editing and comparable neighborhood specificity on ZsRE, with slightly lower generalization, maintaining consistent results across GPT-2 and GPT-J decoders, underscoring its model-agnostic editing capabilities.
@@ 5.3 Sequential Fact Editing
We evaluated Larimar’s ability to perform sequential editing, following the setup of [Hartvigsen et al., 2022], which tackles the issue of forgetting previous edits after multiple sequential edits. Hartvigsen et. al. introduced a continual editing method that integrates an adaptor to update a codebook of edit key-value pairs with a pre-trained language model (GPT2-XL), showing memory retention during sequential edits. We adapt Larimar to this experimental setup, wherein a subset of 200 facts with 5 rephrasings each is selected from the ZsRE validation dataset for testing. In Larimar, a sequential edit is handled by updating the global memory through Eq. (4), again requiring no gradient-based update on incoming edits. For each edit, the encoding of the rephrased query concatenated with the corresponding answer is written to memory. We assessed Larimar’s performance, compared to GRACE, using the edit retention rate (ERR), which is the mean F1 score after 1000 sequential edits when querying the memory with the encoded query $\mathbf{Z}_{\rm query}$ for each written fact. Larimar is not finetuned on question-answer data; instead, we write each question-answer pair as a fact in the memory and query the memory with the original question. We report on the Test Retention Rate (TRR) by evaluating Larimar decoder’s perplexity on 1000 random test samples from wikitext using a separate language model. In comparison, baseline models compute TRR from mean F1 scores from 1000 random samples of NQ data. Results show Larimar’s comparable ERR performance to GRACE, while preserving its original test set performance. Notably, Larimar-1.3B achieves editing speeds approximately 10 or more times faster than GRACE on GPT-2 XL.
Editor Test Retention Rate Edit Retention Rate MEND 0.25 0.27 GRACE 0.69 0.93 Larimar-1.3B 14.6* 0.97 Larimar-6B 15.9* 0.92
Table 3: Sequential editing on ZsRE dataset, showing Larimar does not forget older edits. *We report perplexity on wikitext estimated by a separate language model as Test Retention Rate for Larimar, whereas mean F1 on NQ test set is reported for MEND and GRACE on GPT2-XL [Hartvigsen et al., 2022].
@@
We also evaluated Larimar’s generalization to rephrased prompts, again comparing to GRACE. We use both (i) a dataset of 1000 ZsRE facts, each with 10 variations, divided into edit and holdout sets, and (ii) an edit/holdout dataset with more rephrasings and fewer ($\approx 500$) ZsRE facts. Our analysis, depicted in Figure 3, examines the mean F1 score on the holdout set against the number of memory writes using the edit set, compared to GRACE on the same datasets. [^2] We use GRACE with $\epsilon_{\rm init}=3.0$ to edit block 4 of T5 [Hartvigsen et al., 2022]. As Larimar has no knowledge of upcoming edits, it starts with near-zero F1; in contrast, GRACE has prior knoweldge from training on the edit set. As the sequence of edits grows, Larimar surpasses GRACE’s generalization performance at around 600 edits.
@@
In these experiments, we use $K=1000$, setting the memory size proportional to the number of facts to be written. We also checked an alternative method (see Appendix E) for computing the reading and writing weights, which uses a Gaussian convolution to store each encoding $\mathbf{z}$ in memory location(s) corresponding to the most similar content in a reference memory $\mathbf{M}^{(\rm ref)}$, and which we found to perform better than the pseudoinverse method of [Pham et al., 2021] when there are a relatively small number of rephrasings per fact (see Appendix, Figure 5).

### v2

@@ Abstract
Efficient and accurate updating of knowledge stored in Large Language Models (LLMs) is one of the most pressing research challenges today. This paper presents Larimar - a novel, brain-inspired architecture for enhancing LLMs with a distributed episodic memory. Larimar’s memory allows for dynamic, one-shot updates of knowledge without the need for computationally expensive re-training or fine-tuning. Experimental results on multiple fact editing benchmarks demonstrate that Larimar attains accuracy comparable to most competitive baselines, even in the challenging sequential editing setup, but also excels in speed—yielding speed-ups of 8-10x depending on the base LLM —as well as flexibility due to the proposed architecture being simple, LLM-agnostic, and hence general. We further provide mechanisms for selective fact forgetting, information leakage prevention, and input context length generalization with Larimar and show their effectiveness. Our code is available at https://github.com/IBM/larimar.
@@ 1 Introduction
Table 1: Requirement and capability comparison between Larimar, and two existing editing methods, ROME and GRACE.
@@ 2 Model Details
Background Information: Generative memory networks are a type of associative memory networks that treat memory ‘read’/‘write’ and addressing operations as Bayesian inference where posteriors are updated when a new data episode arrives. Generative Pseudo-inverse Memory (GPM) framework proposed in [Pham et al., 2021] reformulates these Bayesian updates as finding least-square solutions to linear systems, thereby enabling fast and efficient memory operations. One can then generate examples similar to a given input based on memory by sampling from the ‘read’ distribution. In the following sections, we will elaborate the training and inference mechanisms, as we adapt that to train a LLM decoder. In Section 3, we will review the ‘write’, ‘read’ and ‘generate’ operations derived in [Pham et al., 2021] in detail, and formulate additional newly proposed ‘sequential writing’ and ‘forgetting/unlearning’ operations.
@@ 3 Memory operations
       // $\mathbf{Z}$ encoding of the episode to be written to memory (i.e. $\mathbf{Z}=e(\mathbf{X})$) Sample $\xi\sim\mathcal{N}(0,\sigma_{\xi}^{2}I)$
Let $\mathbf{Z}_{\xi}=\mathbf{Z}+\xi$
Compute addressing weight $\mathbf{W}_{0}=\mathbf{Z}_{\xi}\mathbf{M}_{0}^{\dagger}$
1        // $\mathbf{M}_{0}$ is a learned parameter representing prior memory Compute posterior memory $\mathbf{M}=\mathbf{W}_{0}^{\dagger}\mathbf{Z}_{\xi}$
return $\mathbf{M}$
@@
       // $\mathbf{M}$ posterior memory from previous write // $\mathbf{Z}$ encoding of the read input (ie. $\mathbf{Z}=e(\mathbf{X})$) Compute mean addressing weight $\overline{\mathbf{W}}=\mathbf{Z}\mathbf{M}^{\dagger}$
Sample $\mathbf{W}\sim\mathcal{N}(\overline{\mathbf{W}},\sigma_{\mathbf{W}}^{2}I)$
4        // $\sigma_{\mathbf{W}}$ is a learned parameter Compute output latent $\mathbf{Z}_{\text{read}}=\mathbf{WM}$
return $\mathbf{Z}_{\text{read}}$
@@
7       // $\mathbf{M}$ is the posterior memory from a previous write Sample $\mathbf{W}\sim\mathcal{N}(0,I)$
Compute output latent $\mathbf{Z}=\mathbf{WM}$
return $\mathbf{Z}$
@@ 5 Results
Implementation: We employed a BERT large encoder [Devlin et al., 2018] combined with either a GPT2-large [Radford et al., 2019] or a GPTJ-6B decoder and a memory matrix (512x768) for our training experiments, naming the resulting models Larimar-1.3B and Larimar-6B, respectively. Our training data comprised 7.6 million examples constructed by splitting WikiText  [Merity et al., 2016] texts to small chunks of 64 tokens. We used existing pretrained weights from Huggingface to initialize the training. From the $Z_{r}$ (the readout vector), using learned linear projection $W_{M}$, the hidden states are transformed and broadcasted to act as a KV cache across all layers. During the decoder forward pass, this compressed KV cache is used as past KV cache values to generate the memory-controlled output. The hidden size for GPTJ, for example, is 4096 with 28 layers. In testing, the Larimar-1.3B model achieved a perplexity of 14.6, while the Larimar-6B model reached 15.9 on 1,000 random WikiText samples, indicating that adding memory barely affects performance. We trained Larimar-6B models for 10 epochs using Adam optimizer, learning rate 5e-6 and batch size 32. For the Larimar-6B’s training, we used a setup with eight NVIDIA A100-80GB GPUs on a single node, utilizing bfloat16 precision and PyTorch Lightning with the DeepSpeed ZeRO Stage 2 for efficient distributed training.
@@ 5.1 Wall Clock time
Table 1 presents the wall clock time for each editing method across 10 edits, calculated within the EasyEdit framework [Yao et al., 2023] on a single A100 GPU. Results show that Larimar is 4-10x times faster compared to ROME [Meng et al., 2022a] and GRACE [Hartvigsen et al., 2022], two most competitive existing LLM editing baselines. Table 7 further provides a edit-time comparison within other existing baselines, as shown in [Yao et al., 2023], establishing Larimar’s advantage on high-speed editing. Table 1 further lists Larimar’s abilities to handle edits in a training- or tracing- free manner, enabling high-speed editing, handling selective forgetting, and maintain ability to sequential editing setup.
Editor Edit Success Paraphrase Neighborhood S M S M S M GPT-2 XL 22.2 -4.8 24.7 -5.0 78.1 5.0 FT 100.0 98.8 87.9 46.6 40.4 -6.2 FT+L 99.1 91.5 48.7 28.9 70.3 3.5 KN 28.7 -3.4 28.0 -3.3 72.9 3.7 KE 84.3 33.9 75.4 14.6 30.9 -11.0 KE-CF 99.9 97.0 95.8 59.2 6.9 -63.2 MEND 99.1 70.9 65.4 12.2 37.9 -11.6 MEND-CF 100.0 99.2 97.0 65.6 5.5 -69.9 ROME 100.0 97.9 96.4 62.7 75.4 4.2 Larimar-1.3B 100.0 99.7 83.5 50.5 74.7 1.8 Larimar-1.3B + 1 rephrase 100.0 99.6 89.8 62.4 73.3 0.5 Larimar-1.3B + 2 rephrases 100.0 99.6 90.8 63.6 73.3 0.6 GPT-J 16.3 -7.2 18.6 -7.4 83.0 7.3 FT 100.0 99.9 96.6 71.0 10.3 -50.7 FT+L 99.6 95.0 47.9 30.4 78.6 6.8 MEND 97.4 71.5 53.6 11.0 53.9 -6.0 ROME 99.9 99.4 99.1 74.1 78.9 5.2 PROMPT 99.7 80.9 91.0 32.9 37.9 -2.8 IKE (w/ 32 demonstrations) 100.0 91.7 95.2 64.5 77.0 35.2 IKE (w/o paraphrases) 100.0 73.8 83.4 IKE (w/o neighbors) 100.0 99.8 11.5 Larimar-6B 99.6 96.0 88.4 54.7 80.4 4.22 Larimar-6B + 1 rephrase 99.7 95.9 92.9 67.0 79.3 3.5 Larimar-6B + 2 rephrases 99.8 95.7 93.6 67.0 79.2 3.38
Table 2: Single fact editing on CounterFact dataset. Top two best editing methods are highlighted. Larimar uses dynamic memory updates with memory-conditioned decoding and does not require gradient update on edit samples, as opposed to methods needing training (FT, FT+L, MEND) or tracing plus decoder updating (ROME) on edit samples (ROME) or in-context demonstrations (IKE) of (paraphrased) edits and neighboring samples retrieved from a corpus. Though generalization increases when Larimar’s memory is augmented with rephrases at test time.
@@ 5.2 Single Fact editing
We compare the performance of Larimar against a number of recently proposed knowledge editing approaches on the CounterFact dataset [Meng et al., 2022a] designed for testing language models handling of counterfactual edits. It includes 21,919 records to assess if the models can learn new facts rather than simply memorizing the target words. Following other works [Meng et al., 2022a, Zheng et al., 2023], we used the first 2000 samples of this dataset and report the average over single fact editing results for Larimar-1.3B and Larimar-6B in Table 2. The baseline performances are taken from [Meng et al., 2022a, Zheng et al., 2023] (see Related Work and Appendix for details on baseline methods). As opposed to training the LLM on edits, or causally tracing the original fact within LLM and updating the relevant parameters to reflect edit, we leverage Larimar’s one-shot memory update for editing. Wherein, the memory posterior is updated as the edit(s) of interest is written, and then the updated memory is queried. The read-out from the memory then conditions the decoder to output the edit.
@@
As can be seen, when compared to existing editing baselines, Larimar achieves comparable performance in successfully editing new facts, and in the ability to handle neighborhood prompts. For example, when compared with ROME, Larimar performs on par when based on GPT-2 XL, and better when based on GPT-J on editing success and neighborhood specificity, while there remains room to improve generalization. When compared to existing in-context editing approaches (PROMPT and IKE) [Zheng et al., 2023], Larimar does not need multiple in-context demonstrations of the edits and its paraphrases, as well as of neighboring facts, to the decoder, which are retrieved from a corpus. However, as shown in Tables 2 and 8, when Larimar has access to one or two additional paraphrase(s) per fact, by writing it in the memory, the generalization performance increases from 88.4 to 93.6. Note that in this setup the average number of added paraphrase per fact is at most two and we queried the model with a paraphrased prompt unseen by the memory. And, writing latent encodings of paraphrases in memory and conditioning the decoder on the memory is more cost-effective than in-context demonstrations, as the context length does not increase with Larimar’s in-memory mechanisms. Ablation experiments in Appendix shows that a scope detector, either trained on Larimar encodings or encodings from an external LLM, helps with better paraphrase generalization, at the cost of sacrificing the neighborhood specificity. In the absence of a scope detector, the same approach of augmenting memory with two additional rephrases provide an additional 2-3% increase in generalization, irrespective of the dataset (see next paragraph). Throughout the paper, Larimar is configured with a scope detector, unless otherwise mentioned.
We also evaluated Larimar on the ZsRE benchmark [Levy et al., 2017], a QA dataset for relation extraction through reading comprehension, with results displayed in Table 12. Performance scores for GPT-2 XL based baselines are cited from [Meng et al., 2022a], whereas performance of ROME on GPT-J was independently estimated by us. Unlike the CounterFact evaluation, this assessment uses exact match counts for scoring $\mathbb{I}[o^{*}=\text{argmax}_{o}\mathbb{P}[o]]$. Compared to baselines, Larimar demonstrates effective editing and comparable neighborhood specificity on ZsRE, with slightly lower generalization, maintaining consistent results across GPT-2 and GPT-J decoders, underscoring its model-agnostic editing capabilities. We again see that test-time augmentation of the memory with additional paraphrases of the fact boosts generalization from 70.4% to 82.2%, with two rephrases written in memory (Table 13).
@@ 5.3 Sequential Fact Editing
We evaluated Larimar on a sequential editing task, following the setup of [Hartvigsen et al., 2022], which tackles the issue of forgetting previous edits after multiple sequential edits. Hartvigsen et. al. introduced a continual editing method that integrates an adaptor to update a codebook of edit key-value pairs with a pre-trained language model (GPT2-XL), showing memory retention during sequential edits. We adapt Larimar to this experimental setup, wherein a subset of 200 facts with 5 rephrasings each is selected from the ZsRE validation dataset for testing. In Larimar, a sequential edit is handled by updating the global memory through Eq. (4), again requiring no gradient-based update on incoming edits. For each edit, the encoding of the rephrased query concatenated with the corresponding answer is written to memory. We assessed Larimar’s performance, compared to GRACE, using the edit retention rate (ERR), which is the mean F1 score after 1000 sequential edits when querying the memory with the encoded query $\mathbf{Z}_{\rm query}$ for each written fact. Larimar is not finetuned on question-answer data; instead, we write each question-answer pair as a fact in the memory and query the memory with the original question. Results show Larimar’s comparable ERR performance to GRACE, that is the best-performing baseline, while Larimar performing editing approximately 10 or more times faster than GRACE on GPT-2 XL. We found the batch editing method SERAC (referred as DEFER in [Hartvigsen et al., 2022];) to perform worse (ERR=0.31) compared to both Larimar (ERR=0.98) and GRACE (ERR=0.96) on sequential editing on an edit dataset containing duplicates of fact and corresponding rephrasings, though SERAC was trained on fact edits.
MEND GRACE Larimar-1.3B Larimar-6B Edit Retention Rate 0.27 0.93 0.97 0.92
Table 3: Sequential editing on ZsRE, showing Larimar does not forget older edits, baselines are from [Hartvigsen et al., 2022].
@@
We also evaluated Larimar’s generalization to rephrased prompts, again comparing to GRACE. We use (i) a dataset of unique 1000 ZsRE facts, each with 10 variations, divided into edit and holdout sets, and (ii) an edit/holdout dataset with more ($\approx 20$ per fact rephrasings and fewer ($\approx 500$) unique ZsRE facts. Our analysis, depicted in Figure 2 (b), examines the mean F1 score on the holdout set against the number of memory writes using the edit set, compared to GRACE on the same datasets. [^2] We use GRACE with $\epsilon_{\rm init}=3.0$ to edit block 4 of T5 [Hartvigsen et al., 2022]. As Larimar has no knowledge of upcoming edits, it starts with near-zero F1; in contrast, GRACE has prior knoweldge from training on the edit set. As the sequence of edits grows, Larimar surpasses GRACE’s generalization performance at around 600 edits.
@@
In these experiments, we use $K=1000$, setting the memory size proportional to the number of facts to be written. We also checked an alternative method (Appendix E) for computing the reading and writing weights, which uses a Gaussian convolution to store each encoding $\mathbf{z}$ in memory location(s) corresponding to the most similar content in a reference memory $\mathbf{M}^{(\rm ref)}$, and which we found to perform better than the pseudoinverse method of [Pham et al., 2021] when there are a relatively small number of rephrasings per fact (Figure 5).